# Complexity of equal 0-surgeries

### v2

@@ Abstract.
We say that two knots are friends if they share the same 00-surgery. Two friends with different sliceness status would provide a counterexample to the $4$-dimensional smooth Poincaré conjecture. Here we create a census of all friends with small crossing numbers $c$ and tetrahedral complexities $t$, and compute their smooth $4$-genera. In particular, we compute the minimum of $c(K)+c(K^{\prime})$ and of $t(K)+t(K^{\prime})$ among all friends $K$ and $K^{\prime}$. Along the way, we classify all 00-surgeries of prime knots of at most $15$ crossings.
@@
00-surgeries, characterizing slopes, exceptional surgeries, knot traces
@@ 1. Introduction
From a given knot $K$ in the $3$-sphere $S^{3}$ we can construct new $3$- and $4$-manifolds by performing Dehn surgery on $K$ or attaching $2$-handles along $K$. In this article, we are interested in pairs of knots $K$ and $K^{\prime}$ that share the same 00-surgery. To construct the 00-surgery $K(0)$ of a knot $K$ we take the knot exterior $S^{3}\setminus\nu K$ of $K$ and glue a solid torus $V$ to it via a diffeomorphism such that the meridian of $V$ gets identified with the longitude of $K$ that is null-homologous in the knot exterior. We say that two (non-isotopic) knots $K$ and $K^{\prime}$ are friends if their 00-surgeries are orientation-preserving diffeomorphic.
@@
We can also create a $4$-manifold from $K$, its trace $X(K)$, obtained by attaching a $4$-dimensional $2$-handle to $D^{4}$ attached along $K$ with 00-framing. By construction the boundary of $X(K)$ is diffeomorphic to the 00-surgery $K(0)$ of $K$. In fact, it is known that the sliceness status of $K$ is determined by its trace, i.e. if two knots $K$ and $K^{\prime}$ share the same trace, then they are either both slice or both non-slice. On the other hand, it is known that there exist knots with the same traces but different $4$-genus [Pic19]. For an excellent survey about other related results, we refer to [Ten23]. Adapting the above notation, we say that two non-isotopic knots are $4$-dimensional friends if their traces are orientation-preservingly diffeomorphic.
@@
One interest in knots with the same 00-surgeries is that such knots provide a potential way to disprove the smooth $4$-dimensional Poincaré conjecture (see for example [FGMW10, MP23]): If there exist friends $K$ and $K^{\prime}$ such that $K$ is slice (i.e. bounds a smooth disk in $D^{4}$) but $K^{\prime}$ is not slice, then the smooth $4$-dimensional Poincaré conjecture is false. If such a pair of knots exists we might expect it to be simple in some way (or if it is not simple we would expect to not find it at all).
@@
One obvious complexity notion for friends $K$ and $K^{\prime}$ would be the sum of the crossing numbers $c(K)+c(K^{\prime})$. In [MP23] a systematic approach for constructing friends with relatively small crossing numbers was explored. In [MP23, Example 4.10] it was shown that $K12n309$ and $-K14n14254$ are friends and the authors briefly raise the question if these friends minimize $c(K)+c(K^{\prime})$ among all friends $(K,K^{\prime})$. [^1] For knots of at most $15$ crossings we use the DT notation [HTW98]. For knots with crossing numbers in $[16,19]$, we use Burton’s notation [Bur20]. We denote by $-K$ the mirror of $K$. Here (and in the rest of the paper) we only list one chirality of the friends. But of course if $(K,K^{\prime})$ are friends, then $(-K,-K^{\prime})$ are also friends. The construction method in [MP23] is based on RGB links, cf. [MP18]. Another successful method for constructing friends is annulus twisting [Oso06], cf. [AJLO15, AT21]. In fact, Figure 2 of [AT21] shows that $K6a1$ and $7819nh\_$ 78 are friends that yield a smaller sum of crossing numbers than the friends $K12n309$ and $-K14n14254$.
@@
Our main result demonstrates that the above friends $(K6a1,19nh\_78)$ realize the minimum among all friends.
@@
The minimum value of $c(K)+c(K^{\prime})$ among all friends $(K,K^{\prime})$ is $25$. This minimum is uniquely realized by the pair $(K6a1,19nh\_78)$.
Another (arguably more natural) complexity notion for a knot $K$ is the tetrahedral complexity $t(K)$, which is defined to be the minimal number of tetrahedra needed to ideally triangulate the complement of $K$. Hyperbolic knots with tetrahedral complexity at most $9$ have been enumerated by Dunfield [Dun20a]. These knots are called the census knots.
@@
The minimum value of $t(K)+t(K^{\prime})$ among all hyperbolic friends $(K,K^{\prime})$ is $12$. This minimum is uniquely realized by $(m224,-v3093)$. [^2] Here we use the SnapPy census notation of these knots [Dun20a]. The first letter yields the tetrahedral complexity: $m$, $s$, $v$, $t$, $o9$, stand for $5$ (or less), $6$, $7$, $8$, and $9$ tetrahedra, respectively. The census manifold $m224$ is the complement of $-K11n38$, while $v3093$ is the complement of $-16nh\_9$.
Both pairs of minimizers $(K6a1,19nh\_78)$ and $(m224,-v3093)$ are also $4$-dimensional friends. On the other hand, we can also ask for the simplest example of friends that are not $4$-dimensional friends. For the crossing number complexity, we can answer this question, while it remains open for the tetrahedral complexity.
@@
The minimum value of $c(K)+c(K^{\prime})$ among all friends $(K,K^{\prime})$ that are not $4$-dimensional friends is $26$. This minimum is realized by two pairs of friends: $(K12n309,-K14n14254)$ and $(K10n10,-16nh\_17)$. Any other potential minimizer is either $(K11n49,K15n103488)$ or a pair of friends consisting of a $6$-crossing and a $21$-crossing knot.
For the proofs of Theorem 1.1 and 1.2 we use the following strategy. We have an upper bound on these complexity notions coming from the friends $K6a1$ and $7819nh\_$ 78. Since the decision problem for compact $3$-manifolds with boundary and closed $3$-manifolds is solved [Kup19], we can enumerate all pairs of knots up to that complexity and check if their 00-surgeries are diffeomorphic. Here the problem is that the decision problem for $3$-manifolds is only partially implemented and that the enumeration of all these pairs has a high complexity. Nevertheless, we were able to create a census of pairs of knots that share the same 00-surgery that contains all friends $(K,K^{\prime})$ of low complexity. To perform the above strategy we have to use several short-cuts. For our proofs, we are using the verified functions from SnapPy [CDGW] and Regina [BBPea], together with code and data from [ABG+23, BBK+, BK24, Bur20, Dun20a, GAP22, GHMR23, LM23, Swe, Sza, Sag, Thi23, FPS24, BKM24a, BKM24b].
@@
$K6a1$ $s912$ $1$ $7819nh\_$ 78 $[10,11]$ $1$ $C^{\infty}$
$K9n4$ $-m199$ $1$ $-18nh\_23$ $o9\_34801$ $1$ $C^{\infty}$
$K10n10$ $-t12200$ 00 $-16nh\_17$ $-t11532$ 00 not $C^{0}$
$K10n10$ $-t12200$ 00 $[17,28]$ $o9\_43446$ 00 $C^{\infty}$
$K10n13$ $m201$ $1$ $[17,23]$ $o9\_34818$ $1$ $C^{\infty}$
$K11n38$ $-m224$ $1$ $-16nh\_9$ $v3093$ $1$ $C^{\infty}$
$K11n49$ $-v3536$ 00 $K15n103488$ $-v3423$ 00
$K11n49$ $-v3536$ 00 $[17,27]$ $o9\_42735$ 00 $C^{\infty}$
$K11n116$ $-t12748$ 00 $2817nh\_$ 28 $-t12607$ 00 $C^{\infty}$
$K12n121$ $s239$ $1$ $[17,25]$ $-t11441$ $1$
$K12n200$ $t09735$ $1$ $[17,23]$ $t11748$ $1$
$K12n309$ $-v3195$ 00 $-K14n14254$ $v2508$ 00 not $C^{0}$
$K12n318$ $-o9\_40519$ 00 $-18nh\_32$ $-o9\_39433$ 00 not $C^{0}$
$K13n469$ $v2272$ 00 $-K13n469$ $-v2272$ 00 $C^{\infty}$
$K13n572$ $10$ $1$ $-K15n89587$ $[10,11]$ $1$ $C^{\infty}$
$K13n1021$ $-t09900$ $1$ $[17,28]$ $-o9\_34908$ $1$
$K13n1021$ $-t09900$ $1$ $K15n101402$ $-o9\_43876$ $1$
$K13n2527$ $10$ $1$ $-K15n9379$ $t12270$ $1$ $C^{\infty}$
$K13n3158$ $o9\_42515$ 00 $4019nh\_$ 40 $o9\_41909$ 00
$K14n3155$ $10$ $1$ $-K14n3155$ $10$ $1$
$K14n3155$ $10$ $1$ $[16,20]$ $t11462$ $1$
$K14n3155$ $10$ $1$ $[16,20]$ $-t11462$ $1$
$K14n3611$ $o9\_33568$ $1$ $[16,25]$ $-o9\_27542$ $1$ $C^{\infty}$
$K14n5084$ $-o9\_33833$ $1$ $[23,35]$ $-o9\_37547$ $1$ $C^{0}$
$K15n19499$ $o9\_37768$ 00 $K15n153789$ $10$ 00 $C^{\infty}$
$K15n19499$ $o9\_37768$ 00 $[17,26]$ $o9\_31828$ 00 $C^{\infty}$
$K15n94464$ $[10,11]$ $1$ $[17,24]$ $-o9\_40081$ $1$ $C^{\infty}$
$K15n101402$ $-o9\_43876$ $1$ $[17,28]$ $-o9\_34908$ $1$
$K15n103488$ $-v3423$ 00 $[17,27]$ $o9\_42735$ 00
$K15n153789$ $10$ 00 $[17,26]$ $o9\_31828$ 00 $C^{\infty}$
$1716nh\_$ 17 $t11532$ 00 $[17,28]$ $-o9\_43446$ 00 not $C^{0}$
$218nh\_$ 2 $v0595$ $1$ $[19,30]$ $t12120$ $1$
$718nh\_$ 7 $v2869$ $1$ $[16,22]$ $t12388$ $1$ $C^{0}$
$1618nh\_$ 16 $-t10974$ 00 $[17,32]$ $o9\_39967$ 00
$419nh\_$ 4 $t07281$ 00 $[23,42]$ $o9\_34949$ 00 $C^{\infty}$
$419nh\_$ 4 $t07281$ 00 $[17,31]$ $o9\_39806$ 00
$[16,20]$ $t11462$ $1$ $[16,20]$ $-t11462$ $1$ $C^{\infty}$
$[17,24]$ $t11900$ $1$ $[16,28]$ $-o9\_40803$ $1$ $C^{0}$
$[17,31]$ $o9\_22951$ $[0,1]$ $[17,31]$ $-o9\_22951$ $[0,1]$ $C^{\infty}$
$[17,31]$ $o9\_39806$ 00 $[23,42]$ $o9\_34949$ 00
$[16,26]$ $o9\_41058$ $[0,1]$ $[16,26]$ $-o9\_41058$ $[0,1]$ $C^{\infty}$
@@
Theorems 1.1, 1.2 and 1.3 follow directly from Theorem 1.4. In Table 1 we also list the smooth $4$-genus $g_{4}$ of these knots. In particular, it turns out that all friends in our census share the same smooth $4$-genus. (They are either slice, i.e. have smooth $4$-genus 00, or have smooth $4$-genus $1$.) We also remark that all friends from our census share the same $\tau$-invariant from Heegaard Floer homology (which is known to not be a trace invariant) and the same $\nu$-invariant (which is determined by the trace) [HMP21].
@@ 1.1. Code and data
The code and additional data accompanying this paper can be accessed at the GitHub page [AKW24].
@@ Acknowledgements
We are happy to thank Ben Burton and Jonathan Spreer for explaining how to search through the Pachner graph with Regina, Léo Mousseau for help with the $4$-genus computations, Lisa Piccirillo for useful discussions and comments, Marithania Silvero for pointing us to [GHMR23] and [Thi23], and Morwen Thistlethwaite for making his data of the prime $20$ crossing knots available to us.
T.A. is supported by the Research Promotion Program for Acquiring Grants in-Aid for Scientific Research (KAKENHI) in Ritsumeikan University. M.K. was partially supported by the SFB/TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG.

### v3

@@ Abstract.
We say that two knots are friends if they share the same 0-surgery. Two friends with different sliceness status would provide a counterexample to the $4$-dimensional smooth Poincaré conjecture. Here we create a census of all friends with small crossing numbers $c$ and tetrahedral complexities $t$, and compute their smooth $4$-genera. In particular, we compute the minimum of $c(K)+c(K^{\prime})$ and of $t(K)+t(K^{\prime})$ among all friends $K$ and $K^{\prime}$. Along the way, we classify all 0-surgeries of prime knots of at most $15$ crossings.
@@
0-surgeries, characterizing slopes, exceptional surgeries, knot traces
@@ 1. Introduction
From a given knot $K$ in the $3$-sphere $S^{3}$ we can construct new $3$- and $4$-manifolds by performing Dehn surgery on $K$ or attaching $2$-handles along $K$. In this article, we are interested in pairs of knots $K$ and $K^{\prime}$ that share the same 0-surgery. To construct the 0-surgery $K(0)$ of a knot $K$ we take the knot exterior $S^{3}\setminus\nu K$ of $K$ and glue a solid torus $V$ to it via a diffeomorphism such that the meridian of $V$ gets identified with the longitude of $K$ that is null-homologous in the knot exterior. We say that two (non-isotopic) knots $K$ and $K^{\prime}$ are friends if their 0-surgeries are orientation-preserving diffeomorphic.
@@
We can also create a $4$-manifold from $K$, its trace $X(K)$, obtained by attaching a $4$-dimensional 22-handle to $D^{4}$ attached along $K$ with 0-framing. By construction the boundary of $X(K)$ is diffeomorphic to the 0-surgery $K(0)$ of $K$. In fact, it is known that the sliceness status of $K$ is determined by its trace, i.e. if two knots $K$ and $K^{\prime}$ share the same trace, then they are either both slice or both non-slice. On the other hand, it is known that there exist knots with the same traces but different $4$-genus [46]. For an excellent survey about other related results, we refer to [51]. Adapting the above notation, we say that two non-isotopic knots are $4$-dimensional friends if their traces are orientation-preservingly diffeomorphic.
@@
One interest in knots with the same 0-surgeries is that such knots provide a potential way to disprove the smooth $4$-dimensional Poincaré conjecture (see for example [24, 39]): If there exist friends $K$ and $K^{\prime}$ such that $K$ is slice (i.e. bounds a smooth disk in $D^{4}$) but $K^{\prime}$ is not slice, then the smooth $4$-dimensional Poincaré conjecture is false. If such a pair of knots exists we might expect it to be simple in some way.
@@
One obvious complexity notion for friends $K$ and $K^{\prime}$ would be the sum of the crossing numbers $c(K)+c(K^{\prime})$. In [39] a systematic approach for constructing friends with relatively small crossing numbers was explored. In [39, Example 4.10] it was shown that $K12n309$ and K14n14254-K14n14254 are friends and the authors briefly raise the question if these friends minimize $c(K)+c(K^{\prime})$ among all friends $(K,K^{\prime})$. [^1] For knots of at most $15$ crossings we use the DT notation [33]. For knots with crossing numbers in $[16,19]$, we use Burton’s notation [17]. We denote by K-K the mirror of $K$. Here (and in the rest of the paper) we only list one chirality of the friends. But of course if $(K,K^{\prime})$ are friends, then $(-K,-K^{\prime})$ are also friends. The construction method in [39] is based on RBG links, cf. [40]. Another successful method for constructing friends is annulus twisting [43], cf. [1, 4]. In fact, Figure 2 of [4] shows that $K6a1$ and 19nh_7819nh\_78 are friends that yield a smaller sum of crossing numbers than the friends $K12n309$ and K14n14254-K14n14254.
@@
Our main result demonstrates that the above friends (K6a1,19nh_78)(K6a1,19nh\_78) realize the minimum among all friends.
@@
The minimum value of $c(K)+c(K^{\prime})$ among all friends $(K,K^{\prime})$ is $25$. This minimum is uniquely realized by the pair (K6a1,19nh_78)(K6a1,19nh\_78).
Another (arguably more natural) complexity notion for a knot $K$ is the tetrahedral complexity $t(K)$, which is defined to be the minimal number of tetrahedra needed to ideally triangulate the complement of $K$. Hyperbolic knots with tetrahedral complexity at most $9$ have been enumerated by Dunfield [22]. These knots are called the census knots.
@@
The minimum value of $t(K)+t(K^{\prime})$ among all hyperbolic friends $(K,K^{\prime})$ is $12$. This minimum is uniquely realized by (m224,v3093)(m224,-v3093). [^2] Here we use the SnapPy census notation of these knots [22]. The first letter yields the tetrahedral complexity: $m$, $s$, $v$, $t$, $o9$, stand for $5$ (or less), $6$, 77, 88, and $9$ tetrahedra, respectively. The census manifold $m224$ is the complement of K11n38-K11n38, while $v3093$ is the complement of 16nh_9-16nh\_9.
Both pairs of minimizers (K6a1,19nh_78)(K6a1,19nh\_78) and (m224,v3093)(m224,-v3093) are also $4$-dimensional friends. On the other hand, we can also ask for the simplest example of friends that are not $4$-dimensional friends. For the crossing number complexity, we can answer this question, while it remains open for the tetrahedral complexity.
@@
The minimum value of $c(K)+c(K^{\prime})$ among all friends $(K,K^{\prime})$ that are not $4$-dimensional friends is $26$. This minimum is realized by two pairs of friends: (K12n309,K14n14254)(K12n309,-K14n14254) and (K10n10,16nh_17)(K10n10,-16nh\_17). Any other potential minimizer is either $(K11n49,K15n103488)$ or a pair of friends consisting of a $6$-crossing and a $21$-crossing knot.
For the proofs of Theorem 1.1 and 1.2 we use the following strategy. We have an upper bound on these complexity notions coming from the friends $K6a1$ and 19nh_7819nh\_78. Since the decision problem for compact $3$-manifolds with boundary and closed $3$-manifolds is solved [36], we can enumerate all pairs of knots up to that complexity and check if their 0-surgeries are diffeomorphic. Here the problem is that the decision problem for $3$-manifolds is only partially implemented and that the enumeration of all these pairs has a high complexity. Nevertheless, we were able to create a census of pairs of knots that share the same 0-surgery that contains all friends $(K,K^{\prime})$ of low complexity. To perform the above strategy we have to use several short-cuts. For our proofs, we are using the verified functions from SnapPy [19] and Regina [18], together with code and data from [7, 8, 12, 17, 22, 28, 30, 38, 49, 50, 48, 52, 26, 10, 9].
@@
$K6a1$ $s912$ 11 19nh_7819nh\_78 $[10,11]$ 11 $C^{\infty}$
$K9n4$ m199-m199 11 18nh_23-18nh\_23 o9_34801o9\_34801 11 $C^{\infty}$
$K10n10$ t12200-t12200 0 16nh_17-16nh\_17 t11532-t11532 0 not $C^{0}$
$K10n10$ t12200-t12200 0 $[17,28]$ o9_43446o9\_43446 0 $C^{\infty}$
$K10n13$ $m201$ 11 $[17,23]$ o9_34818o9\_34818 11 $C^{\infty}$
$K11n38$ m224-m224 11 16nh_9-16nh\_9 $v3093$ 11 $C^{\infty}$
$K11n49$ v3536-v3536 0 $K15n103488$ v3423-v3423 0
$K11n49$ v3536-v3536 0 $[17,27]$ o9_42735o9\_42735 0 $C^{\infty}$
$K11n116$ t12748-t12748 0 17nh_2817nh\_28 t12607-t12607 0 $C^{\infty}$
$K12n121$ $s239$ 11 $[17,25]$ t11441-t11441 11
$K12n200$ $t09735$ 11 $[17,23]$ $t11748$ 11
$K12n309$ v3195-v3195 0 K14n14254-K14n14254 $v2508$ 0 not $C^{0}$
$K12n318$ o9_40519-o9\_40519 0 18nh_32-18nh\_32 o9_39433-o9\_39433 0 not $C^{0}$
$K13n469$ $v2272$ 0 K13n469-K13n469 v2272-v2272 0 $C^{\infty}$
$K13n572$ $10$ 11 K15n89587-K15n89587 $[10,11]$ 11 $C^{\infty}$
$K13n1021$ t09900-t09900 11 $[17,28]$ o9_34908-o9\_34908 11
$K13n1021$ t09900-t09900 11 $K15n101402$ o9_43876-o9\_43876 11
$K13n2527$ $10$ 11 K15n9379-K15n9379 $t12270$ 11 $C^{\infty}$
$K13n3158$ o9_42515o9\_42515 0 19nh_4019nh\_40 o9_41909o9\_41909 0
$K14n3155$ $10$ 11 K14n3155-K14n3155 $10$ 11
$K14n3155$ $10$ 11 $[16,20]$ $t11462$ 11
$K14n3155$ $10$ 11 $[16,20]$ t11462-t11462 11
$K14n3611$ o9_33568o9\_33568 11 $[16,25]$ o9_27542-o9\_27542 11 $C^{\infty}$
$K14n5084$ o9_33833-o9\_33833 11 $[23,35]$ o9_37547-o9\_37547 11 $C^{0}$
$K15n19499$ o9_37768o9\_37768 0 $K15n153789$ $10$ 0 $C^{\infty}$
$K15n19499$ o9_37768o9\_37768 0 $[17,26]$ o9_31828o9\_31828 0 $C^{\infty}$
$K15n94464$ $[10,11]$ 11 $[17,24]$ o9_40081-o9\_40081 11 $C^{\infty}$
$K15n101402$ o9_43876-o9\_43876 11 $[17,28]$ o9_34908-o9\_34908 11
$K15n103488$ v3423-v3423 0 $[17,27]$ o9_42735o9\_42735 0
$K15n153789$ $10$ 0 $[17,26]$ o9_31828o9\_31828 0 $C^{\infty}$
16nh_1716nh\_17 $t11532$ 0 $[17,28]$ o9_43446-o9\_43446 0 not $C^{0}$
18nh_218nh\_2 $v0595$ 11 $[19,30]$ $t12120$ 11
18nh_718nh\_7 $v2869$ 11 $[16,22]$ $t12388$ 11 $C^{0}$
18nh_1618nh\_16 t10974-t10974 0 $[17,32]$ o9_39967o9\_39967 0
19nh_419nh\_4 $t07281$ 0 $[23,42]$ o9_34949o9\_34949 0 $C^{\infty}$
19nh_419nh\_4 $t07281$ 0 $[17,31]$ o9_39806o9\_39806 0
$[16,20]$ $t11462$ 11 $[16,20]$ t11462-t11462 11 $C^{\infty}$
$[17,24]$ $t11900$ 11 $[16,28]$ o9_40803-o9\_40803 11 $C^{0}$
$[17,31]$ o9_22951o9\_22951 $[0,1]$ $[17,31]$ o9_22951-o9\_22951 $[0,1]$ $C^{\infty}$
$[17,31]$ o9_39806o9\_39806 0 $[23,42]$ o9_34949o9\_34949 0
$[16,26]$ o9_41058o9\_41058 $[0,1]$ $[16,26]$ o9_41058-o9\_41058 $[0,1]$ $C^{\infty}$
Theorems 1.1, 1.2 and 1.3 follow directly from Theorem 1.4. In Table 1 we also list the smooth $4$-genus $g_{4}$ of these knots. In particular, it turns out that all friends in our census share the same smooth $4$-genus. (They are either slice, i.e. have smooth $4$-genus 0, or have smooth $4$-genus 11.) In particular, this rules out a large class of potential candidates for pairs of knots with the same 0-surgery that could be used to construct a counterexample to the smooth $4$-dimensional Poincaré conjecture. If a pair of friends leading to a counterexample to the smooth $4$-dimensional Poincaré conjecture exists, at least one of them needs to have large crossing number (at least $16$) and tetrahedral complexity (at least $10$). Thus, a brute force approach like that done in this article is not very likely to find such a pair of friends. Another computational strategy to search for a pair of friends with different sliceness statuses is explained in [21] and studied further in [20].
@@
We also remark that all friends from our census share the same $\tau$-invariant from Heegaard Floer homology (which is known to not be a trace invariant) and the same $\nu$-invariant (which is determined by the trace) [31].
@@ 1.1. Code and data
The code and additional data accompanying this paper can be accessed at the GitHub page [3].
@@ Acknowledgements
We are happy to thank Ben Burton and Jonathan Spreer for explaining how to search through the Pachner graph with Regina, Léo Mousseau for help with the $4$-genus computations, Lisa Piccirillo for useful discussions and comments, Marithania Silvero for pointing us to [30] and [52], and Morwen Thistlethwaite for making his data of the prime $20$ crossing knots available to us. Moreover, we thank the referee for careful proofreading.
Individual grant support
T.A. is supported by the Research Promotion Program for Acquiring Grants in-Aid for Scientific Research (KAKENHI) in Ritsumeikan University.
MK is supported by the DFG, German Research Foundation, (Project: 561898308) and by the SFB/TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics; by a Ramón y Cajal grant (RYC2023-043251-I) and the project PID2024-157173NB-I00 funded by MCIN/AEI/10.13039/501100011033, by ESF+, and by FEDER, EU; and by a VII Plan Propio de Investigación y Transferencia (SOL2025-36103) of the University of Sevilla.